OpenStudy (dqswag):

What is the Value of g(-6) g(x)= -3 + 4|x|

hartnn (hartnn):

what have you tried? did you plug in x=-6 in g(x)?

OpenStudy (dqswag):

What is the Value of g(-6) g(x)= -3 + 4|x| g(-6)=

OpenStudy (dqswag):

Will I have to multiply 6 to 4 than subtract by 3??

hartnn (hartnn):

to get g(-6), plug in x = -6 in g(x) -3 + 4|-6| do you know what is the value of |-6|?

OpenStudy (dqswag):

yes it will be 24 right

hartnn (hartnn):

you mean |-6| will be 6 and 4*6 = 24. ?

hartnn (hartnn):

-3 + 4|-6| = -3 4*6 = -3+24 = 21 :)

OpenStudy (dqswag):

Thx bra!

hartnn (hartnn):

welcome ^_^